# Highest-weight vectors and three-point functions in GKO coset decomposition

### v1

@@ 4.4 Virasoro conformal blocks
$\mathrm{F}_{b}(\vec{P},\vec{P^{\prime}};\vec{z})=\langle v_{P_{n},b},\Phi_{\Delta_{n-1}}(z_{n-1})\dots\Phi_{\Delta_{3}}(z_{3})\Phi_{\Delta_{2}}(z_{2}),v_{P_{1},b}\rangle,$ (4.31)
@@
Let $\Phi_{\Delta_{2}}(z)\colon\mathbb{M}_{P_{1},b}\rightarrow\overline{\mathbb{M}_{P_{3},b}}$, $\Delta_{2}=\Delta(P_{2},b)$ be a Virasoro vertex operator as in Sec. 2.7. The vector $\mathbb{W}^{b;P_{2}}_{P_{1},P_{3}}(z)=\Phi_{\Delta}(z)|v_{P_{1},b}\rangle\in\overline{\mathbb{M}_{P_{3},b}}$ is called Virasoro chain vector.
It follows from statements in Sec. 2.7 that for generic values of parameters the chain vector $\mathbb{W}^{b;P_{2}}_{P_{1},P_{3}}(z)$ exists and uniquely fixed by its top component, i.e. by scalar product $\langle v_{P_{3},b},\mathbb{W}^{b;P_{2}}_{P_{1},P_{3}}(z)\rangle$. For conformal blocks we have
@@
$\left(z(1-z)\partial_{z}^{2}+b^{2}\Big{(}(2z-1)\partial_{z}+\frac{\Delta_{1}}{z}+\frac{\Delta_{3}}{z-1}+\Delta_{12}-\Delta_{4}\Big{)}\right)\mathrm{F}(z)=0$ (4.33)
@@ 4.5 $\widehat{\mathfrak{sl}}(2)$ conformal blocks
Assume that there are given parameters $\lambda_{i}$, $i\in\{1,\dots,n\}$ and $\lambda_{j}^{\prime}$, $j\in\{2,\dots,n-2\}$. Assume also that there are given vertex operators $\mathcal{V}_{\lambda_{2}}(x,z)\colon\mathcal{M}_{\lambda_{1},k}\rightarrow\overline{\mathcal{M}_{\lambda^{\prime}_{2},k}}$, $\mathcal{V}_{\lambda_{i}}(z)\colon\mathcal{M}_{\lambda^{\prime}_{i-1},k}\rightarrow\overline{\mathcal{M}_{\lambda^{\prime}_{i},k}}$ for $i\in\{3,\dots,n-2\}$, and $\mathcal{V}_{\lambda_{n-1}}(z)\colon\mathcal{M}_{\lambda^{\prime}_{n-2},k}\rightarrow\overline{\mathcal{M}_{\lambda_{n},k}}$ as in Sec. 2.6. The $n$-point sphere conformal block for $\widehat{\mathfrak{sl}}(2)$ algebra defined as
$\Psi_{k}(\vec{\lambda},\vec{\lambda^{\prime}};\vec{x},\vec{z})=\langle v_{\lambda_{n},k},\mathcal{V}_{\lambda_{n-1}}(x_{n-1},z_{n-1})\dots\mathcal{V}_{\lambda_{3}}(x_{3},z_{3})\mathcal{V}_{\lambda_{2}}(x_{2},z_{2}),v_{\lambda_{1},k}\rangle,$ (4.39)
@@
Let $\mathcal{V}_{\lambda_{2}}(x,z)\colon\mathcal{M}_{\lambda_{1},k}\rightarrow\mathcal{M}_{\lambda_{3},k}$ be an $\widehat{\mathfrak{sl}}(2)$ vertex operator as in Sec. 2.6. Vector $\mathcal{W}^{k;\lambda_{2}}_{\lambda_{1},\lambda_{3}}(x,z)=\mathcal{V}_{\lambda_{2}}(x,z)|v_{\lambda_{1},k}\rangle\in\overline{\mathcal{M}_{\lambda_{3},k}}$ is called $\widehat{\mathfrak{sl}}(2)$-chain vector.
Consider four-point conformal blocks. Similarly to Virasoro case one can assume that $z_{2}=z$, $x_{2}=x$, $z_{3}=1$, $z_{3}=1$ and we have only one intermediate parameter $\lambda_{2}^{\prime}=\lambda^{\prime}$. We have
$\Psi_{k}(\vec{\lambda},\lambda^{\prime};x,z)=\langle v_{\lambda_{4},k},\mathcal{V}_{\lambda_{3}}(1,1)\mathcal{V}_{\lambda_{2}}(x,z),v_{\lambda_{1},k}\rangle=\langle\mathcal{W}^{k;\lambda_{3}}_{\lambda_{4},\lambda^{\prime}}(-1,1),\mathcal{W}^{k;\lambda_{2}}_{\lambda_{1},\lambda^{\prime}}(x,z)\rangle.$ (4.40)
Unless otherwise stated, we assume that Whittaker vectors are normalized by $\langle v_{\lambda_{3},k},\mathcal{W}^{k;\lambda_{2}}_{\lambda_{1},\lambda_{3}}(x,z)\rangle=z^{\Delta_{3}-\Delta_{1}-\Delta_{2}}x^{(\lambda_{1}+\lambda_{2}-\lambda_{3})/2}$. Hence conformal blocks have the form
$\Psi_{k}(\vec{\lambda},\lambda^{\prime};x,z)=z^{\Delta^{\prime}-\Delta_{1}-\Delta_{2}}x^{(\lambda_{1}+\lambda_{2}-\lambda^{\prime})/2}(1+O(x)+O(z)).$ (4.41)
We will need description of four-point conformal block with insertion of degenerate field of spin 1 which we denote by $X_{1}(x,z)$, see (2.46). This description follows from Knizhnik-Zamolodchikov equations [KZ84] and is well known, see e.g. [Tes99]. Consider function
@@
$\kappa\partial_{z}\Psi_{k,1}$ $=\frac{\partial_{y}\Psi_{k,0}-\frac{1}{2}(-2y\partial_{y}+\nu)\Psi_{k,1}}{z-1}+\frac{\frac{1}{2}(-\lambda-2)\Psi_{1}-\partial_{y}\Psi_{k,0}}{z}.$ (4.43b)
@@ 4.6 Coset decomposition and conformal block relations
Now, we are ready to come back to Coset Construction. The vector $\widetilde{\mathcal{U}}_{\nu,n}(x,z)|\tilde{u}_{l}(\lambda)\rangle$ would be a tensor product of Virasoro and $\widehat{\mathfrak{sl}}(2)$-chain vectors since the operator $\widetilde{\mathcal{U}}_{\nu,n}$ is primary fiels for $\mathrm{Vir}^{\mathrm{coset}}\oplus\widehat{\mathfrak{sl}}(2)_{k+1}$. A chain vector in each summand of (3.1) would be a tensor product of Virasoro and $\widehat{\mathfrak{sl}}(2)$-chain vectors. We fix normalization of such products by
$\langle\mathbb{W}^{b;P(\nu)+nb}_{P(\lambda)+lb,P(\mu)+mb}(z)\otimes\mathcal{W}^{k+1,\nu+2n}_{\lambda+2l,\mu+2m,}(x,z)|\tilde{u}_{m}(\mu)\rangle=\|\tilde{u}_{m}(\mu)\|^{2}z^{\Delta(P(\mu))-\Delta(P(\nu))-\Delta(P(\lambda))}x^{\frac{-\mu+\nu+\lambda}{2}-m+n+l}.$ (4.48)
@@
$\widetilde{\mathcal{U}}_{\nu,n}(x,z)|\tilde{u}_{l}(\lambda)\rangle$ $=\sum_{n\in\mathbb{Z}}\frac{\tilde{C}_{m,n,l}(\mu,\nu,\lambda)}{\|\tilde{u}_{m}(\mu)\|^{2}}\mathbb{W}^{b;P(\nu)+nb}_{P(\lambda)+lb,P(\mu)+mb}(z)\otimes\mathcal{W}^{k+1,\nu+2n}_{\lambda+2l,\mu+2m}(x,z),$ (4.49)
$b(x,z)|\tilde{u}_{l}(\lambda)\rangle$ $=\sum_{s\in\pm 1}\frac{\tilde{B}^{s}_{l}(\lambda)}{\|\tilde{u}_{l+{s}/{2}}(\lambda)\|^{2}}\mathbb{W}^{b,P_{1,2}(b)}_{P(\lambda)+lb,P(\lambda)+l+\frac{s}{2}}(z)\otimes\mathcal{W}^{k+1;1}_{\lambda+2l,\lambda+2l+s}(x,z),$ (4.50)
@@ 4.6.1 Recurrence relations
$\langle\tilde{u}_{m}(\mu)|\widetilde{\mathcal{U}}_{\nu,n}(1,1)b(x,z)|\tilde{u}_{l}(\lambda)\rangle.$ (4.51)
Due to Proposition 4.23 it decomposes as a linear combination of Virasoro and $\widehat{\mathfrak{sl}}$ conformal blocks. Due to Proposition 4.12 these are conformal blocks with the presence of degenerate fields. Therefore we get
@@
The four-point conformal block with presence of fields $b(x,z)$ have the form
$\langle\tilde{u}_{m}(\mu),\widetilde{\mathcal{U}}_{\nu,n}(1,1)b(x,z)\,\tilde{u}_{l}(\lambda)\rangle=\sum_{s\in\pm 1}\frac{\tilde{C}_{m,n,l+s/2}(\mu,\nu,\lambda)\tilde{B}^{s}_{l}(\lambda)}{\|\tilde{u}_{l+s/2}(\lambda)\|^{2}}\\ \times\mathrm{F}_{b}\left(\vec{P}_{\vec{l}},P(\lambda)+lb+\frac{s}{2}b;z\right)\Psi_{k+1}\left(\vec{\lambda}_{\vec{l}},\lambda+2l+s;x,z\right)$ (4.52)
@@
$\mathrm{F}_{b}\left(\vec{P}_{\vec{l}},P(\lambda)+lb-b/2b;z\right)\Psi_{k+1}\left(\vec{\lambda}_{\vec{l}},\lambda+2l-1;x,z\right)=(1-z)^{\frac{1}{2}(-r_{1}-r_{2}+r_{3})}z^{\frac{1+r_{3}}{2}}\\ \Big{(}\frac{a_{2}-a_{3}}{a_{3}-1}z\,{}_{2}F_{1}(\hat{I}\vec{a}|z)\,{}_{2}F_{1}(\hat{I}(-\vec{a}-\vec{r})|z)+x\,{}_{2}F_{1}(\hat{I}(\vec{a}+(1,0,1))|z)\,{}_{2}F_{1}(\hat{I}(-\vec{a}-\vec{r})|z)\Big{)},$ (4.54b)
@@
$a_{2}$ $=\frac{(\lambda+2l+\mu+2m+\nu+2n+3)}{2(\kappa+1)},$ $~{}~{}r_{2}$ $=-l-m-n-\frac{1}{2};$ (4.55b)
@@
On the other hand, it follows from the definition of $b(x,z)$ in formula (2.51) that the action $b(x,z)$ on $\tilde{u}_{l}(\lambda)$ is given by Laurent series in $z$ for $l\in\mathbb{Z}$ and by Laurent series in $z$ times $z^{1/2}$ for $l\in\mathbb{Z}+1/2$. Similarly singular part of OPE of $b(x,z)$ and $\widetilde{\mathcal{U}}_{\nu,n}(1,1)$ is either Laurent polynomial in $z-1$ if $n\in\mathbb{Z}$ or Laurent polynomial in $z-1$ times $(z-1)^{1/2}$ for $n\in\mathbb{Z}+1/2$. In any case, after factoring out term $(1-z)^{\frac{1}{2}(-r_{1}-r_{2}+r_{3})}z^{\frac{-1-r_{3}}{2}}$ we get linear combination of products of ${}_{2}F_{1}$ hypergeometric functions which has trivial monodromy in $z$ and can have only poles for $z=0,1,\infty$. Hence it is a rational function in $z$. This gives strong restriction on coefficients in this linear combination. Namely
@@
where $p_{\vec{r}}(\vec{a})$ is a coefficient found in Proposition A.2. Using Corollary 3.17 and Proposition 4.13 after straightforward computation we get
@@ 4.6.2 Conformal block relations and blowup equations
$\langle\tilde{u}_{0}(\mu_{4})|\widetilde{\mathcal{U}}_{\mu_{3},0}(1,1)\widetilde{\mathcal{U}}_{\mu_{2},0}(x,z)|\tilde{u}_{0}(\mu_{1})\rangle.$ (4.58)
Since the $\widehat{\mathfrak{sl}}(2)_{1}$ part of the vectors $u_{0}(\mu_{4}),u_{0}(\mu_{1})$ and vertex operators $\widetilde{\mathcal{U}}_{\mu_{3},0},\widetilde{\mathcal{U}}_{\mu_{2},0}$ are trivial. Then this conformal block is equal to $\Psi_{k}\left(\vec{\mu},\lambda;x,z\right)$. On the other hand we can use decomposition (3.1) and write
@@
$\vec{P}=(P(\mu_{1}),P(\mu_{2}),P(\mu_{3}),P(\mu_{4})),\quad\vec{\mu}=(\mu_{1},\mu_{2},\mu_{3}.\mu_{4}).$ (4.60)
Due to AGT correspondence the function $\mathrm{F}_{b}$ is equal (up to simple $U(1)$ factor) to Nekrasov partition function for $SU(2)$ theory with $N_{f}=4$ [AGT10] and the function $\Psi_{k}$ is equal to Nekrasov partition function for $SU(2)$ with presence of the surface defect [AT10], [Nek17], [NT22]. In this geometric language the relation (4.59) is blowup relation with the presence of a surface defect, suggested in [Nek24], [JN20] (equations without defect were proven in [NY05]).
@@
$\Psi_{k}^{\text{tor}}(\lambda,\lambda^{\prime};x,q)=\operatorname{Tr}\left(q^{L_{0}}x^{h_{0}}\mathcal{V}_{\lambda}(1,1)\right)|_{\mathcal{M}_{\lambda^{\prime},k}},\quad\mathrm{F}_{b}^{\text{tor}}(P^{\prime},P;q)=\operatorname{Tr}\left(q^{L_{0}}\Phi_{\Delta}(z)\right)|_{\mathbb{M}_{P^{\prime},b}}.$ (4.61)
@@
Clearly one can generalize such relation for more point conformal blocks on a sphere or torus. In the next section, we will also consider Whittaker limit of these relations.

### v3

@@ 4.4 Virasoro conformal blocks
$\mathrm{F}_{b}(\vec{P},\vec{P^{\prime}};\vec{z})=\Big{\langle}v_{P_{n},b},\Phi_{\Delta_{n-1}}(z_{n-1})\dots\Phi_{\Delta_{3}}(z_{3})\Phi_{\Delta_{2}}(z_{2})v_{P_{1},b}\Big{\rangle},$ (4.31)
@@
Let $\Phi_{\Delta_{2}}(z)\colon\mathbb{M}_{P_{1},b}\rightarrow\overline{\mathbb{M}_{P_{3},b}}$, $\Delta_{2}=\Delta(P_{2},b)$ be a Virasoro vertex operator as in sec. 2.7. The vector $\mathbb{W}^{b;P_{2}}_{P_{1},P_{3}}(z)=\Phi_{\Delta}(z)v_{P_{1},b}\in\overline{\mathbb{M}_{P_{3},b}}$ is called Virasoro chain vector.
It follows from statements in sec. 2.7 that for generic values of parameters the chain vector $\mathbb{W}^{b;P_{2}}_{P_{1},P_{3}}(z)$ exists and uniquely fixed by its top component, i.e. by scalar product $\big{\langle}v_{P_{3},b},\mathbb{W}^{b;P_{2}}_{P_{1},P_{3}}(z)\big{\rangle}$. For conformal blocks we have
@@
$\left(z(1-z)\partial_{z}^{2}+b^{2}\Big{(}(2z-1)\partial_{z}+\frac{\Delta_{1}}{z}+\frac{\Delta_{3}}{z-1}+\Delta_{12}-\Delta_{4}\Big{)}\right)\mathrm{F}(z)=0.$ (4.33)
@@ 4.5 $\widehat{\mathfrak{sl}}(2)$ conformal blocks
Assume that there are given parameters $\lambda_{i}$, $i\in\{1,\dots,n\}$ and $\lambda_{j}^{\prime}$, $j\in\{2,\dots,n-2\}$. Assume also that there are given vertex operators $\mathcal{V}_{\lambda_{2}}(x,z)\colon\mathcal{M}_{\lambda_{1},k}\rightarrow\overline{\mathcal{M}_{\lambda^{\prime}_{2},k}}$, $\mathcal{V}_{\lambda_{i}}(x,z)\colon\mathcal{M}_{\lambda^{\prime}_{i-1},k}\rightarrow\overline{\mathcal{M}_{\lambda^{\prime}_{i},k}}$ for $i\in\{3,\dots,n-2\}$, and $\mathcal{V}_{\lambda_{n-1}}(x,z)\colon\mathcal{M}_{\lambda^{\prime}_{n-2},k}\rightarrow\overline{\mathcal{M}_{\lambda_{n},k}}$ as in sec. 2.6. The $n$-point sphere conformal block for $\widehat{\mathfrak{sl}}(2)$ algebra defined as
$\Psi_{k}(\vec{\lambda},\vec{\lambda^{\prime}};\vec{x},\vec{z})=\Big{\langle}v_{\lambda_{n},k},\mathcal{V}_{\lambda_{n-1}}(x_{n-1},z_{n-1})\dots\mathcal{V}_{\lambda_{3}}(x_{3},z_{3})\mathcal{V}_{\lambda_{2}}(x_{2},z_{2})v_{\lambda_{1},k}\Big{\rangle},$ (4.39)
@@
Let $\mathcal{V}_{\lambda_{2}}(x,z)\colon\mathcal{M}_{\lambda_{1},k}\rightarrow\mathcal{M}_{\lambda_{3},k}$ be an $\widehat{\mathfrak{sl}}(2)$ vertex operator as in sec. 2.6. Vector $\mathcal{W}^{k;\lambda_{2}}_{\lambda_{1},\lambda_{3}}(x,z)=\mathcal{V}_{\lambda_{2}}(x,z)v_{\lambda_{1},k}\in\overline{\mathcal{M}_{\lambda_{3},k}}$ is called $\widehat{\mathfrak{sl}}(2)$-chain vector.
Consider four-point conformal blocks. Similarly to Virasoro case one can assume that $z_{2}=z$, $x_{2}=x$, $z_{3}=1$, $x_{3}=1$ and we have only one intermediate parameter $\lambda_{2}^{\prime}=\lambda^{\prime}$. We have
$\Psi_{k}(\vec{\lambda},\lambda^{\prime};x,z)=\Big{\langle}v_{\lambda_{4},k},\mathcal{V}_{\lambda_{3}}(1,1)\mathcal{V}_{\lambda_{2}}(x,z)v_{\lambda_{1},k}\Big{\rangle}=\Big{\langle}\mathcal{W}^{k;\lambda_{3}}_{\lambda_{4},\lambda^{\prime}}(-1,1),\mathcal{W}^{k;\lambda_{2}}_{\lambda_{1},\lambda^{\prime}}(x,z)\Big{\rangle}.$ (4.40)
Unless otherwise stated, we assume that Whittaker vectors are normalized by $\big{\langle}v_{\lambda_{3},k},\mathcal{W}^{k;\lambda_{2}}_{\lambda_{1},\lambda_{3}}(x,z)\big{\rangle}=z^{\Delta_{3}-\Delta_{1}-\Delta_{2}}x^{(\lambda_{1}+\lambda_{2}-\lambda_{3})/2}$. Hence conformal blocks have the form
$\Psi_{k}(\vec{\lambda},\lambda^{\prime};x,z)=z^{\Delta^{\prime}-\Delta_{1}-\Delta_{2}}x^{(\lambda_{1}+\lambda_{2}-\lambda^{\prime})/2}\Big{(}1+O(x)+O(z)\Big{)}.$ (4.41)
We will need a description of four-point conformal block with insertion of degenerate field of spin $\frac{1}{2}$ which we denote by $X_{1}(x,z)$, see (2.46). This description follows from Knizhnik-Zamolodchikov equations [KZ84] and is well known, see e.g. [Tes99]. Consider function
@@
$\kappa\partial_{z}\Psi_{k,1}$ $=\frac{\partial_{y}\Psi_{k,0}-\frac{1}{2}(-2y\partial_{y}+\nu)\Psi_{k,1}}{z-1}+\frac{\frac{1}{2}(-\lambda-2)\Psi_{k,1}-\partial_{y}\Psi_{k,0}}{z}.$ (4.43b)
@@ 4.6 Coset decomposition and conformal block relations
Now, we are ready to come back to coset construction. The vector $\widetilde{\mathcal{U}}_{\nu,n}(x,z)\tilde{u}_{l}(\lambda)$ would be a tensor product of Virasoro and $\widehat{\mathfrak{sl}}(2)$-chain vectors since the operator $\widetilde{\mathcal{U}}_{\nu,n}$ is a primary field for $\mathrm{Vir}^{\mathrm{coset}}\oplus\widehat{\mathfrak{sl}}(2)_{k+1}$. A chain vector in each summand of (3.1) would be a tensor product of Virasoro and $\widehat{\mathfrak{sl}}(2)$-chain vectors. We fix normalization of such products by
$\Big{\langle}\mathbb{W}^{b;P(\nu)+nb}_{P(\lambda)+lb,P(\mu)+mb}(z)\otimes\mathcal{W}^{k+1,\nu+2n}_{\lambda+2l,\mu+2m}(x,z),\tilde{u}_{m}(\mu)\Big{\rangle}\\ =\|\tilde{u}_{m}(\mu)\|^{2}z^{\Delta(P(\mu))-\Delta(P(\nu))-\Delta(P(\lambda))}x^{\frac{-\mu+\nu+\lambda}{2}-m+n+l}.$ (4.48)
@@
$\widetilde{\mathcal{U}}_{\nu,n}(x,z)\tilde{u}_{l}(\lambda)$ $=\sum_{n\in\mathbb{Z}}\frac{\tilde{C}_{m,n,l}(\mu,\nu,\lambda)}{\|\tilde{u}_{m}(\mu)\|^{2}}\,\mathbb{W}^{b;P(\nu)+nb}_{P(\lambda)+lb,P(\mu)+mb}(z)\otimes\mathcal{W}^{k+1,\nu+2n}_{\lambda+2l,\mu+2m}(x,z),$ (4.49)
$b(x,z)\tilde{u}_{l}(\lambda)$ $=\sum_{s\in\pm 1}\frac{\tilde{B}^{s}_{l}(\lambda)}{\|\tilde{u}_{l+{s}/{2}}(\lambda)\|^{2}}\,\mathbb{W}^{b,P_{1,2}(b)}_{P(\lambda)+lb,P(\lambda)+l+\frac{s}{2}}(z)\otimes\mathcal{W}^{k+1;1}_{\lambda+2l,\lambda+2l+s}(x,z),$ (4.50)
@@ 4.6.1 Recurrence relations
$\Big{\langle}\tilde{u}_{m}(\mu),\widetilde{\mathcal{U}}_{\nu,n}(1,1)\,b(x,z)\,\tilde{u}_{l}(\lambda)\Big{\rangle}.$ (4.51)
Due to Proposition 4.23, it decomposes as a linear combination of Virasoro and $\widehat{\mathfrak{sl}}(2)$ conformal blocks. According to Proposition 4.12, these are conformal blocks with the presence of degenerate fields. Therefore we get
@@
The four-point conformal block with presence of fields $b(x,z)$ has the form
$\langle\tilde{u}_{m}(\mu),\widetilde{\mathcal{U}}_{\nu,n}(1,1)b(x,z)\,\tilde{u}_{l}(\lambda)\rangle=\sum_{s\in\pm 1}\frac{\tilde{C}_{m,n,l+s/2}(\mu,\nu,\lambda)\tilde{B}^{s}_{l}(\lambda)}{\|\tilde{u}_{l+s/2}(\lambda)\|^{2}}\\ \times\mathrm{F}_{b}\left(\vec{P}_{\vec{l}},P(\lambda)+lb+sb/2;z\right)\Psi_{k+1}\left(\vec{\lambda}_{\vec{l}},\lambda+2l+s;x,z\right),$ (4.52)
@@
$\mathrm{F}_{b}\left(\vec{P}_{\vec{l}},P(\lambda)+lb-b/2;z\right)\Psi_{k+1}\left(\vec{\lambda}_{\vec{l}},\lambda+2l-1;x,z\right)=(1-z)^{\frac{1}{2}(-r_{1}-r_{2}+r_{3})}z^{\frac{1+r_{3}}{2}}\\ \Big{(}\frac{a_{2}-a_{3}}{a_{3}-1}z\,{}_{2}F_{1}(\hat{I}\vec{a}|z)\,{}_{2}F_{1}(\hat{I}(-\vec{a}-\vec{r})|z)+x\,{}_{2}F_{1}(\hat{I}(\vec{a}+(1,0,1))|z)\,{}_{2}F_{1}(\hat{I}(-\vec{a}-\vec{r})|z)\Big{)},$ (4.54b)
@@
$a_{2}$ $=\frac{(\lambda+2l+\mu+2m+\nu+2n+3)}{2(\kappa+1)},$ $~{}~{}r_{2}$ $=-l-m-n-\frac{1}{2},$ (4.55b)
@@
On the other hand, it follows from the definition of $b(x,z)$ in formula (2.51) that the action $b(x,z)$ on $\tilde{u}_{l}(\lambda)$ is given by Laurent series in $z$ for $l\in\mathbb{Z}$ and by Laurent series in $z$ times $z^{1/2}$ for $l\in\mathbb{Z}+1/2$. Similarly singular part of OPE of $b(x,z)$ and $\widetilde{\mathcal{U}}_{\nu,n}(1,1)$ is either Laurent polynomial in $z-1$ if $n\in\mathbb{Z}$ or Laurent polynomial in $z-1$ times $(z-1)^{1/2}$ for $n\in\mathbb{Z}+1/2$. In any case, after factoring out the term $(1-z)^{\frac{1}{2}(-r_{1}-r_{2}+r_{3})}z^{\frac{-1-r_{3}}{2}}$ we get a linear combination of products of ${}_{2}F_{1}$ hypergeometric functions which has trivial monodromy in $z$ and can have only poles at $z=0,1,\infty$. Hence it is a rational function in $z$. This gives strong restriction on the coefficients in this linear combination. Namely
@@
where $p_{\vec{r}}(\vec{a})$ is a coefficient found in Proposition A.2. Using Corollary 3.18 and Proposition 4.13 after straightforward computation we get a recursion formula.
@@ 4.6.2 Conformal block relations and blowup equations
$\Big{\langle}\tilde{u}_{0}(\mu_{4}),\widetilde{\mathcal{U}}_{\mu_{3},0}(1,1)\,\widetilde{\mathcal{U}}_{\mu_{2},0}(x,z)\,\tilde{u}_{0}(\mu_{1})\Big{\rangle}.$ (4.58)
Note that the $\widehat{\mathfrak{sl}}(2)_{1}$ part of the vectors $u_{0}(\mu_{4}),u_{0}(\mu_{1})$ and vertex operators $\widetilde{\mathcal{U}}_{\mu_{3},0},\widetilde{\mathcal{U}}_{\mu_{2},0}$ are trivial. Hence this conformal block is equal to $\Psi_{k}\left(\vec{\mu},\lambda;x,z\right)$. On the other hand, we can use decomposition (3.1) and write
@@
$\vec{P}=(P(\mu_{1}),P(\mu_{2}),P(\mu_{3}),P(\mu_{4})),\quad\vec{\mu}=(\mu_{1},\mu_{2},\mu_{3},\mu_{4}).$ (4.60)
Due to AGT correspondence the function $\mathrm{F}_{b}$ is equal (up to simple $U(1)$ factor) to Nekrasov partition function for $SU(2)$ theory with $N_{f}=4$ [AGT10] and the function $\Psi_{k}$ is equal to Nekrasov partition function for $SU(2)$ theory with presence of a surface defect [AT10], [Nek17], [NT22]. In this geometric language the relation (4.59) is a blowup relation with the presence of the surface defect, suggested in [Nek24], [JN20] (equations without defect were proven by Nakajima and Yoshioka in the seminal paper [NY05]).
@@
$\Psi_{k}^{\text{tor}}(\lambda,\lambda^{\prime};x,q)=\operatorname{Tr}\left(q^{L_{0}}x^{h_{0}}\mathcal{V}_{\lambda}(1,1)\right)|_{\mathcal{M}_{\lambda^{\prime},k}},\quad\mathrm{F}_{b}^{\text{tor}}(P,P^{\prime};q)=\operatorname{Tr}\left(q^{L_{0}}\Phi_{\Delta}(z)\right)|_{\mathbb{M}_{P^{\prime},b}}.$ (4.61)
@@
Clearly one can generalize such relations for more point conformal blocks on sphere or torus. In the next section, we will also consider Whittaker limit of these relations.